OpenStudy (anonymous):

Find the area of the trapezoid A)95 units2 B)190 units2 C)150 units2 D)300 units2

OpenStudy (anonymous):

helpp

OpenStudy (anonymous):

\[Area = \left( \frac{ b1 + b2 }{ 2 } \right) h\]

OpenStudy (anonymous):

what is the answer

OpenStudy (anonymous):

115

OpenStudy (anonymous):

ther is no 155

OpenStudy (anonymous):

the answers r at the top

OpenStudy (anonymous):

\[A = \left( \frac{ 11 + 12}{ 2 } \right) 10 = 115\]

OpenStudy (anonymous):

ther is no 115 r u sure that right

OpenStudy (anonymous):

Are you sure that's the right trapezoid and the base and side lengths are correct?

OpenStudy (anonymous):

ya it right

OpenStudy (anonymous):

I'm positive for that trapezoid the answer is 115

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):